### v1

# Photonic Doping of Epsilon-Near-Zero Bragg Microcavities

Abstract
Epsilon-near-zero (ENZ) photonics provides a unique route to extreme dispersion, strong field confinement, and unconventional wave phenomena. Among its most intriguing concepts is photonic doping, where subwavelength dielectric inclusions embedded in ENZ media unlock exotic responses such as perfect magnetic conductor behavior and simultaneous epsilon-and-mu-near-zero properties. While prior demonstrations of photonic doping have been limited to microwave and far-infrared domains due to material losses, we demonstrate the feasibility of photonic doping at optical frequencies. Specifically, we implement photonic doping by embedding nanoparticle Mie resonators into an ultra-low-loss, all-dielectric ENZ platform realized with near-cutoff Bragg-reflection microcavities. We reveal the emergence of hybrid Bragg–Mie, near-zero-index modes whose fields localize within or between the nanoparticles. These modes exhibit substantially higher quality (Q) and Purcell factors than either the standalone nanoparticle array or the bare cavity. This mechanism enables the formation of highly isolated, spectrally pure electric and magnetic modes with dipolar or multipolar character. In representative photonically doped Bragg cavities, we predict $Q$-factors on the order of $10^{4}$ and magnetic-dipole Purcell factors exceeding $4\times 10^{3}$ in the near-infrared region, associated with strong and high-Q magnetic hot spots localized within the nanoparticles. These results establish photonic doping of ENZ Bragg cavities as a powerful paradigm for achieving ultra-narrow-bandwidth, pure electric or magnetic Mie resonances, with exceptional potential for low-threshold nonlinear optics, magnetic dipole spectroscopy in atomic systems, and strong quantum optical light–matter interactions.
keywords:
American Chemical Society,
[…]
{tocentry}
Some journals require a graphical entry for the Table of Contents. This should be laid out “print ready” so that the sizing of the text is correct.
Inside the tocentry environment, the font used is Helvetica 8 pt, as required by Journal of the American Chemical Society.
The surrounding frame is 9 cm by 3.5 cm, which is the maximum permitted for Journal of the American Chemical Society graphical table of content entries. The box will not resize if the content is too big: instead it will overflow the edge of the box.
This box and the associated title will always be printed on a separate page at the end of the document.
1 1. Introduction
Controlling light at the nanoscale lies at the heart of nanophotonics and relies fundamentally on the interaction between light and matter. According to Maxwell’s equations, the optical response of a medium is governed by its electric permittivity ($\varepsilon$) and magnetic permeability ($\mu$), which together define the refractive index $n$. Traditionally, photonic engineering has focused on spatial and temporal modulation of these parameters, using refractive index contrast and dispersion to manipulate the propagation of light.
Recent advances in material science and nanofabrication have enabled unprecedented spatiotemporal control over $\varepsilon$, $\mu$, and $n$, giving rise to new regimes of light–matter interaction. Among the most intriguing is the field of near-zero-index (NZI) photonics 1, encompassing materials in which the effective refractive index approaches zero due to vanishing $\varepsilon$, $\mu$, or both, commonly referred to as epsilon-near-zero (ENZ), mu-near-zero (MNZ), and epsilon-and-mu-near-zero (EMNZ) media. In NZI regimes, the optical wavelength becomes effectively infinite, the phase velocity diverges, and the wavevector tends toward zero. However, some aspects of light–matter interaction, such as group velocity, wave impedance, and field confinement, vary significantly across NZI classes 2, leading to distinct features and functionalities which underpin effects like perfect transmission, cloaking, super-coupling, diffraction suppression and enhanced light–matter interactions 3.
A recent approach to tune or modify the electromagnetic behavior of ENZ media has been termed photonic doping, by analogy with semiconductor doping, as subwavelength dielectric inclusions impart entirely new functionalities to ENZ materials. 4. This strategy has enabled extraordinary phenomena, including tunneling through ENZ media or general impedance matching 5, 6, 7, geometry-independent radiation of antennas 8, dispersion coding 9, magnetic field concentration or enhancement 10 and finally, realization of perfect magnetic conductor (PMC) and EMNZ behavior by applying non-magnetic components 4, 5. However, material losses and fabrication challenges at optical wavelengths have so far limited the photonic doping of ENZ media primarily to microwave and far-infrared frequencies.
In this work, we present the first demonstration of photonic doping of ENZ media in the optical domain by employing all-dielectric Bragg-reflection (BR) microcavities engineered to operate near their cutoff frequencies, thereby emulating ENZ behavior at optical wavelengths 11, 12. We investigate the general behavior of spherical and cylindrical inclusions in ENZ media and focus specifically on photonic doping using priodic array of cylindrical dielectric nanocylinders (NCs) embedded within ENZ BR microcavities. Our results show that this configuration can yield EMNZ and generate NZI hybrid Bragg-Mie modes with significantly enhanced quality ($Q$) and Purcell factors (PFs), surpassing those of the bare cavity and the isolated NC array. This enhancement arises from improved field confinement and reduced mode volume induced by the embedded NCs, leading to emergence of high $Q$-factor magnetic hot-spots in analogy with electric hot-spots in plasmonic nanostructures.
We identify two distinct classes of NZI modes in the doped cavity. The first, features strong $Q$-factor enhancement and is characterized by fields predominantly confined within the cavity core, either inside or between the NCs. The second class exhibits fields largely distributed within the Bragg mirror layers and achieves ultra-high Q-factors ($\sim 10^{5}$) nearly one order of magnitude higher than those of the core-confined modes. The huge $Q$-factor enhancement of Mie resonances compared to individual or coupled NC arrays, results in highly isolated and spectrally pure electric or magnetic modes, with dipolar or multipolar character.
Mie theory offers an exact analytical solution to Maxwell’s equations for light scattering from highly symmetric particles such as spheres and cylinders. It describes the emergence of intrinsic electric and magnetic multipolar eigenmodes, known as Mie resonances, when the incident wavelength matches conditions set by the particle’s size, geometry, and refractive index contrast with its surroundings 13. These quantities are conveniently unified in the so-called size parameter, which combines wavelength, refractive index, and particle size. Here, we demonstrate that in photonically doped BR cavities, the incidence angle of the illuminating wave and the cutoff wavelength of the Bragg cavity provide additional degrees of freedom, beyond geometry and material properties, for spectral tuning of the Mie-like resonances.
Although electric and magnetic Mie resonances with dipolar or multipolar character typically occur at different frequencies, they often spectrally overlap, making it challenging to isolate pure modal responses 14. Photonically doped Bragg cavities provide a unique route to achieving pure magnetic dipole scattering, which is crucial for unambiguous studies of magnetic light–matter interactions in nonmagnetic nanoparticles (NPs) and for the selective excitation of intrinsic magnetic dipole transitions in materials 15, 16. In analogy to tailoring the electric local density of states (LDOS) for controlling electric dipole transitions, the magnetic LDOS can be engineered to enable or enhance magnetic dipole processes 17, 18. Previous efforts based on anapole modes in engineered nanostructures, including core–shell nanospheres 19, high-index dielectrics 15, and metal–dielectric hybridized nanodisks 20, have demonstrated suppression of electric dipole radiation. However, these mechanisms remain limited in purity and tunability compared to photonic doping in Bragg cavities, which yields highly isolated and extremely high-$Q$ magnetic dipole resonances.
Light–matter interactions at optical frequencies, are typically dominated by the electric field component of light, since the magnetic response of most materials is weak. Nevertheless, achieving strong magnetic light–matter coupling is essential for a range of applications, including Purcell enhancement of magnetic dipole transitions in materials and quantum emitters 21, 15, 22, 23, development of optical metamaterials 24, 25, 26, and the enhancement of nonlinear optical effects 27, 28, 29, 30.
In what follows, Section 2 introduces an analytical study of Mie resonances in ENZ media, independent of the specific mechanism by which the ENZ condition is realized. Section 3 reports a numerical investigation of photonic doping in ENZ Bragg cavities, addressing the excitation of both transverse electric and transverse magnetic resonances, the tunability of hybrid Bragg–Mie modes, and the enhancement of the magnetic PF. Finally, Section 4 summarizes the main findings and discusses their implications.
2. Analytical Study of Mie Resonances in ENZ Media
[…]
In this section, we provide an analytical demonstration of this effect for spherical and cylindrical dielectric objects, based on Mie scattering theory. We begin by applying Mie theory to an infinitely long dielectric cylinder of radius $R$, illuminated by a normally incident plane wave with wavenumber $k$. For the case where the electric field is polarized perpendicular to the cylinder axis (TE polarization), the extinction efficiency $Q_{\mathrm{ext}}$ is proportional to the real part of the scattering coefficients 13, given by
$a_{n}=\frac{x_{i}J_{n}(x_{i})J_{n}^{\prime}(x_{e})-x_{e}J_{n}(x_{e})J_{n}^{\prime}(x_{i})}{x_{i}J_{n}(x_{i}){H_{n}^{(1)}}^{\prime}(x_{e})-x_{e}H_{n}^{(1)}(x_{e})J_{n}^{\prime}(x_{i})},$ (1)
where, $J_{n}$ and $H_{n}^{(1)}$ denote the Bessel and Hankel functions of the first kind, respectively. The size parameters are defined as $x_{i}=n_{i}kR$ and $x_{e}=n_{e}kR$, corresponding to the arguments of the Bessel and Hankel functions inside and outside the cylinder. In these expressions, $n_{i}$ and $n_{e}$ denote the refractive indices of the cylinder and the surrounding medium, $k$ is the vacuum wavenumber, and $R$ is the cylinder radius.
In Fig. 1a, the real parts of the Mie coefficients $a_{0}$, $a_{1}$, and $a_{2}$ are plotted as functions of wavelength, representing the electric and magnetic resonances of an infinite dielectric cylinder with a radius $R=150~\text{nm}$. The cylinder ($n_{i}=3.5$) is embedded in a medium with $n_{e}=1$ and is illuminated by a normally incident plane wave whose electric field is perpendicular to the cylinder axis (see inset).
Figure 1: Real parts of Mie coefficients $a_{0}$, $a_{1}$, and $a_{2}$ as functions of wavelength for an infinite cylinder of radius $R$, plotted using the paramters R=150R=150~nm, $n_{i}=3.5$ and (a) $n_{e}=1$, (b) $n_{e}=0.1$. The inset illustrates normal incidence of a plane wave with its electric field polarized perpendicular to the cylinder axis.
Fig. 1b presents the same Mie coefficients, with identical parameters except for the refractive index of the surrounding medium, which is set to $n_{e}=0.1$. This modification results in a pronounced narrowing of the resonances. In particular, for the TE0 magnetic dipole mode (corresponding to $a_{0}$ Mie coefficient), the bandwidth decreases from approximately 200 nanometers down to around $2\,\mathrm{nm}$ at $n_{e}=0.1$. We adopt a dispersionless, purely real index $n_{i}=3.5$ over the entire spectral range of $\lambda=200$$2000~\mathrm{nm}$ to decouple geometry-driven modal physics from material-specific line-shaping. In this idealized setting, electric and magnetic dipolar and higher-order (angular) modes, together with their radial overtones, exhibit narrowed bandwidths in the ENZ regime. Material dispersion and loss would chiefly add spectral shifts and broadening, masking–but not altering–the underlying mechanism.
This resonance narrowing in the ENZ regime, where $n_{e}\to 0$ and $x_{e}\ll 1$, can be analytically justified by applying Eq. (1) and using the small-argument approximations for the functions $J_{n}(x_{e})$, $J_{n}^{\prime}(x_{e})$, $H_{n}^{(1)}(x_{e})$, and ${H_{n}^{(1)}}^{\prime}(x_{e})$. As demonstrated in Appendix A, this procedure yields the resonance condition of $J_{n}(x_{i})=0$ (consistent with the results shown in Fig. 1b), while for the resonance width ($\gamma_{n}$) of the TEn modes, we obtain
$\gamma_{n}\propto\frac{1}{2(n+1)!}\left[x_{e}^{2}J_{n}^{\prime}(x_{i})\right]\left(\frac{x_{e}}{2}\right)^{n+1},$ (2)
which vanishes in the limit $x_{e}\to 0$.
For the case of TM modes, where the electric field is polarized along the cylinder axis (shown in the inset of Fig. 2a), the extinction efficiency is proportional to the real part of the scattering coefficients $b_{n}$, expressed as:
$b_{n}=\frac{x_{e}J_{n}(x_{i})J_{n}^{\prime}(x_{e})-x_{i}J_{n}(x_{e})J_{n}^{\prime}(x_{i})}{x_{e}J_{n}(x_{i}){H_{n}^{(1)}}^{\prime}(x_{e})-x_{i}H_{n}^{(1)}(x_{e})J_{n}^{\prime}(x_{i})}.$ (3)
The real parts of the Mie coefficients $b_{0}$, $b_{1}$, and $b_{2}$ are plotted in Fig. 2, for the above-considered photonically doped system. Significant narrowing of the resonances is clear upon comparison of Figs. 2a and 2b, corresponding to $n_{e}=1$ and $n_{e}=0.1$.
Figure 2: Real parts of Mie coefficients $b_{0}$, $b_{1}$, and $b_{2}$ as functions of wavelength using the same parameters as in Fig. 1 and for (a) $n_{e}=1$, (b) $n_{e}=0.1$.
By using similar small-argument approximations for Bessel and Hankel functions, it can be shown (Appendix A) that the resonance condition for TMn modes with $n>0$ is satisfied when:
$\frac{n}{2}J_{n}(x_{i})+x_{i}J_{n}^{\prime}(x_{i})=0.$ (4)
However, as discussed in Appendix A, the TM0 mode which corresponds to a cylindrically symmetric electric monopole field distribution, does not support a true resonance, as can also be inferred from the red curve in Fig. 2b. For the resonance width of the TMn modes with $n>0$ we get:
$\gamma_{n}\propto\frac{1}{n!}\left[\frac{n}{2}J_{n}(x_{i})-x_{i}J_{n}^{\prime}(x_{i})\right]\left(\frac{x_{e}}{2}\right)^{n},$ (5)
which vanishes in the ENZ regime as $x_{e}\to 0$, indicating strong suppression of radiative losses and enhanced confinement.
The $Q$-factor enhancement and slight blue shift of Mie resonances in the ENZ regime, as demonstrated in Figs. 1 and 2, can be attributed to the extremely high impedance contrast between the dielectric NC and its ENZ background. This contrast suppresses scattering or radiative losses, narrowing the resonance bandwidth, while simultaneously inhibiting field penetration into the surrounding medium. The resulting strong field confinement reduces the effective mode volume, making the particle behave like a smaller optical cavity and shifting the resonance to shorter wavelengths.
Similar effects can be observed for a spherical particle of radius $R$, characterized by the size parameters $x_{i}=n_{i}kR$ and $x_{e}=n_{e}kR$, corresponding to the interior and exterior of the sphere, respectively. When the sphere is illuminated with a plane wave with wavenumber $k$, the Mie coefficients for the electric ($a_{n}$) and magnetic ($b_{n}$) multipole modes can be written in terms of the spherical Bessel functions $j_{n}(x)$ and $h_{n}^{(1)}(x)$ or equivalently the Riccati–Bessel functions $\psi_{n}(x)=xj_{n}(x)$ and $\xi_{n}(x)=xh_{n}^{(1)}(x)$ as:
$a_{n}=\frac{x_{i}\psi_{n}(x_{i})\psi_{n}^{\prime}(x_{e})-x_{e}\psi_{n}(x_{e})\psi_{n}^{\prime}(x_{i})}{x_{i}\psi_{n}(x_{i})\xi_{n}^{\prime}(x_{e})-x_{e}\xi_{n}(x_{e})\psi_{n}^{\prime}(x_{i})},$ (6)
$b_{n}=\frac{x_{e}\psi_{n}(x_{i})\psi_{n}^{\prime}(x_{e})-x_{i}\psi_{n}(x_{e})\psi_{n}^{\prime}(x_{i})}{x_{e}\psi_{n}(x_{i})\xi_{n}^{\prime}(x_{e})-x_{i}\xi_{n}(x_{e})\psi_{n}^{\prime}(x_{i})}.$ (7)
In Appendix B, it is shown analytically that in the ENZ regime, as $x_{e}\to 0$, the $Q$-factors of all electric and magnetic dipolar and higher-order resonances are enhanced. To illustrate this effect, the real parts of Mie coefficients, corresponding to electric dipole ($a_{1}$), electric quadrupole ($a_{2}$), magnetic dipole ($b_{1}$) and magnetic quadrupole ($b_{2}$) resonances of a spherical NP of radius R=85R=85~nm and refractive index $n_{i}=3.4$ are plotted for two cases of $n_{e}=1$ and $n_{e}=0.1$, shown in Figs. 3a and 3b, respectively.
Figure 3: Real parts of Mie coefficients as functions of wavelength for a spherical NP of radius R=85nmR=85~nm and refractive index $n_{i}=3.4$, inside a medium with refractive index of (a) $n_{e}=1$, (b) $n_{e}=0.1$.
These results show that, in the $\varepsilon\to 0$ limit, the electric dipolar and higher-order Mie resonances of spherical dielectric nanoparticles are also enhanced and spectrally narrowed, thereby extending the earlier prediction of Ref. 31, which emphasized this effect for magnetic dipolar and multipolar resonances.
3. Numerical Study of Photonic Doping in all-dielectric ENZ Cavities
As noted earlier, most studies of photonic doping in ENZ media have examined near-cutoff metallic or PEC waveguides emulating ENZ behavior. Remarkably, inserting even a single dielectric inclusion into such structures can drastically alter their effective permeability, giving rise to PMC- or EMNZ-like responses. In contrast, these effects do not clearly manifest when a single NP is embedded in the core of an ENZ BR cavity. To observe similar phenomena in this configuration, an array of NPs is required. The contrast comes from the way the NPs reshape the BR cavity mode. A sparse NP population, i.e. a large-period lattice, with the extreme case of a single NP, drives the hybrid mode to delocalize from the core and leak into the Bragg stacks, degrading confinement and thus reducing the modal overlap between the Mie resonance and the cavity ENZ mode. In the opposite limit, a denser array (smaller period) supports collective Mie behavior and an effective–medium response that pulls the field back into the core. This restores cavity–Mie modal overlap, yielding sharper resonances (higher $Q$) and stronger coupling, with energy concentrated in the core rather than dissipated in the mirrors.
To numerically investigate these effects and the photonic doping features of ENZ cavities, we examine the resonance behavior of an array of infinitely long dielectric cylinders embedded in ENZ BR cavities. Two-dimensional Comsol simulations are performed to calculate the transmittance of the structure shown in Fig. 4, and the eigenmodes are obtained using the mode analysis module.
Figure 4: Simulation domain of the ENZ Bragg cavity loaded with an array of infinitely-long dielectric cylinders of radius $R=150~\mathrm{nm}$ and refractive index $n=3.5$.
1.1 3.1. Excitation of Transverse Electric (TE) Resonances
Figure 5 presents the transmittance spectra of a cavity, both with and without the NC array in the core. The cavity consists of a half-wave SiO2 core with refractive index of $n_{\mathrm{SiO_{2}}}=1.44+i10^{-6}$ 32, sandwiched between 10 pairs of quarter-wave cladding layers of SiO2 and SiN ($n_{\mathrm{SiN}}=2+i10^{-7}$)33. The cutoff wavelength of the cavity is set to $\lambda_{c}=1500~\mathrm{nm}$. We dope the system with an array of silicon NCs with radii $R=150~\mathrm{nm}$ and refractive indices $n_{\mathrm{Si}}=3.5+i1.38\times 10^{-13}$ 34, positioned at the center of the core. The periodicity of the array is $p=500~\mathrm{nm}$ and $p=700~\mathrm{nm}$ in Figs. 5b and 5c, respectively.
Figure 5: Transmittance spectra and corresponding field profiles: (a) bare Bragg cavity (cutoff wavelength $\lambda_{c}=1500\,\mathrm{nm}$) consisting a half-wave SiO2 core, sandwiched between 10 pairs of quarter-wave cladding layers of SiO2 and SiN; (b,c) the same cavity loaded with arrays of infinitely-long silicon NCs, placed at the core center, with lattice periods (b) $p=500\,\mathrm{nm}$ and (c) $p=700\,\mathrm{nm}$. All structures are illuminated by a normally incident plane wave (along $y$-axis) with the magnetic field $H$ along $z$-axis (TE, $H_{z}$).
Figure 5a shows the transmittance spectrum and field profiles of the bare cavity mode, with a transmission peak at $\lambda=1500~\mathrm{nm}$. The electric field is mainly concentrated at the core center, while the magnetic field is localized at the core boundaries. Introducing the NC array (Fig. 5b) gives rise to two resonant modes within the Bragg stop band, one at a longer wavelength and the other at a shorter wavelength relative to the bare-cavity cutoff. As demonstrated later, these correspond to NZI modes of the structure. The higher-energy mode, located at $\lambda=1371~\mathrm{nm}$, is dominantly confined inside the cylinders and can be identified as the Mie TE0-like mode, associated with the magnetic dipole resonance of the Mie coefficient $a_{0}$. This mode exhibits an enhanced quality factor of $Q=626$, nearly $50$ times higher than its free-space resonance. The lower-energy NZI mode, at $\lambda=1623.5~\mathrm{nm}$ with a quality factor of $Q=607$, is characterized by distinct electric and magnetic field distributions that are mainly localized outside the cylinders. Increasing the periodicity of the array preserves the general field distribution patterns of both the high- and low-energy modes. However, the $Q$-factors of the modes increase, and as shown in Fig. 5c, the lower-energy mode exhibits a blue shift while the higher-energy mode undergoes a red shift, which can extend beyond the cutoff wavelength of the bare cavity (see Fig. 6).
By increasing the periodicity to $p=900~\mathrm{nm}$, the quality factor of the TE0-like mode rises to $Q=2.32\times 10^{4}$, which is about $2\times 10^{3}$ times higher than its free-space Mie counterpart and more than 14 times higher than that of the bare cavity. As shown in Fig.6, additional resonant structural modes also appear, with electromagnetic fields predominantly distributed within the Bragg layers but exhibiting exceptionally high quality factors; for example, $Q=4.87\times 10^{5}$ at $\lambda=1459.1~\mathrm{nm}$, more than 21 times higher than the confined TE0-like mode.
Figure 6: Transmittance spectrum and field profiles of the resonant modes in the same Bragg cavity of Fig. 5b, but with the period of $p=900~\mathrm{nm}$, illuminated by an incident TE plane wave with the magnetic field $H$ normal to the image plane.
Comparison of the magnetic field distribution of the TE0-like mode with that of the bare-cavity mode (Fig. 5a) reveals that enhanced field localization, together with the significant increase in $Q$-factor, produces a magnetic hot spot within the NC, which is one of the key features of photonic doping in ENZ media 10.
Figure 7: Transmittance spectra of a single NC of radius $R=150~\mathrm{nm}$ (solid curve) in a SiO2 matrix as well as arrays of such NCs with periods varying between 400-900 nm.
In these structures, the periodicity of the NC array plays a crucial role in achieving high-$Q$ resonances. However, the effect of periodicity on the transmission peak of the same array in free space is opposite. To demonstrate this, we calculated the transmittance of a plane wave for a single NC and for arrays of identical NCs with varying periodicity. The results show that arranging the NCs into an array causes the transmission peak to broaden and shift to shorter wavelengths, with the effect becoming more pronounced at smaller periods.
1.2 3.1.1 PMC and EMNZ Features in Photonically Doped BR Cavities
To demonstrate the PMC and EMNZ features by photonic doping of BR cavities, the same structure as in Fig. 5b is considered, but with increased number of Bragg layers to $N=14$ pairs of quarter-wave layers. The transmittance spectrum of the structure is shown in Fig. 8a with the corresponding field profiles of the TE0-like mode at $\lambda=1372.38~\mathrm{nm}$. We have also calculated the dispersive effective refractive index (using the mode analysis module of COMSOL), as well as the effective relative permittivity and permeability around this mode. The results are shown in Fig. 8b.
Figure 8: (a) Transmittance and resonant-mode field distribution corresponding to a BR cavity with $N=14$ pairs of cladding layers on each side of the half-wave core and doped with an array of cylindrical NCs of radius $R=150~\mathrm{nm}$ having a period of $p=500~\mathrm{nm}$. (b) Calculated effective refractive index as well as effective relative permittivity and permeability around the wavelength $\lambda=1372.38~\mathrm{nm}$ corresponding to the TE0-like resonant mode in the same NC-doped BR cavity.
In order to consistently define the effective magnetic permeability of a periodic array of inclusions, we follow the homogenization procedure introduced by Silveirinha and Engheta 35. For a periodic Bloch mode, the microscopic electric and polarization current density fields have the form 36
$\mathbf{E}(\mathbf{r})=\tilde{\mathbf{E}}(\mathbf{r})\,e^{i\mathbf{k}_{B}\cdot\mathbf{r}},\qquad\mathbf{J}(\mathbf{r})=\tilde{\mathbf{J}}(\mathbf{r})\,e^{i\mathbf{k}_{B}\cdot\mathbf{r}},$ (8)
where $\tilde{\mathbf{E}}$ and $\tilde{\mathbf{J}}$ are periodic functions over the unit cell and $\mathbf{k}_{B}$ is the Bloch wavevector. Accordingly, the microscopic magnetic dipole moment per unit length is written as 36
$m_{z}\;=\;\frac{1}{2}\iint_{\text{inc}}\Big[(x-x_{c})J_{y}-(y-y_{c})J_{x}\Big]\,e^{-i\mathbf{k}_{B}\cdot\mathbf{r}}\,dA,$ (9)
where $J_{x}=-i\omega(\varepsilon-\varepsilon_{\text{host}})E_{x}$ and $J_{y}=-i\omega(\varepsilon-\varepsilon_{\text{host}})E_{y}$ are the polarization current densities inside the inclusion, and $(x_{c},y_{c})$ is the centroid of the inclusion cross-section, included to avoid origin dependence. The corresponding magnetization is normalized over the whole unit cell area,
$M_{z}=\frac{m_{z}}{A_{\text{cell}}},\qquad A_{\text{cell}}=\iint_{\text{cell}}dA.$ (10)
The dephasing factor $e^{-i\mathbf{k}_{B}\cdot\mathbf{r}}$ is also used consistently in the evaluation of all other cell-averaged quantities, such as $\langle H_{z}\rangle$, ensuring that the retrieved effective permeability $\mu_{\text{eff}}$ is invariant under translations of the unit cell and fully consistent with Floquet–Bloch homogenization theory. However, since the mode under consideration is an NZI mode, with effective index close to zero, the Bloch wavevector is negligible and the dephasing factor can be safely ignored.
Since the direct cell average of the magnetic field, $\langle H_{z}\rangle$, contains contributions from both the applied field and the self-field of the induced dipole, a corrected bulk magnetic field is introduced as 5
$H_{\text{bulk}}=\langle H_{z}\rangle-\frac{m_{z}}{A_{\text{cell}}}.$ (11)
The constitutive relation can then be written as
$B=\mu_{0}\langle H_{z}\rangle=\mu_{\text{eff}}H_{\text{bulk}},$ (12)
which leads to the following expression for the effective relative permeability:
$\mu_{r}=\frac{\mu_{\text{eff}}}{\mu_{0}}=\frac{\langle H_{z}\rangle}{H_{\text{bulk}}}.$ (13)
In Fig. 8b, the real and imaginary parts of the effective index of refraction are shown by the dark-blue and red-circle curves, respectively. The real and imaginary parts of the effective relative permeability are plotted as green curves, while the effective relative permittivity, calculated from the relation $\varepsilon_{\mathrm{eff}}=n_{\mathrm{eff}}^{2}/\mu_{\mathrm{eff}}$, is shown by the dotted light-blue curve.
The TE0-like mode resonates at $\lambda=1372.38~\mathrm{nm}$, in the close vicinity of the cutoff. The PMC-like response appears at the wavelength where the effective permeability (solid green) exhibits a pole, $\mu_{r}\to\infty$, corresponding to a high-impedance condition. The EMNZ region, corresponding to wavelengths shorter than the cutoff wavelength, is highlighted in blue. In this region, both the effective permittivity and permeability are positive and smaller than unity. For wavelengths longer than the cutoff wavelength, the signs of the effective permittivity and permeability are opposite, and consequently the wave cannot propagate in this spectral range, as indicated by the rapid growth of the imaginary part of the effective index (red-circle curve).
It is interesting to note that, although the microscopic magnetic dipole moment $m_{z}$ and consequently the TE0-like mode exhibits a resonance at $\lambda=1372.38~\mathrm{nm}$, the effective relative permeability diverges at a slightly longer wavelength. This red shift follows directly from the retrieval formula
$\mu_{r}=\frac{\langle H_{z}\rangle}{\langle H_{z}\rangle-M_{z}},\qquad M_{z}=\frac{m_{z}}{A_{\text{cell}}}.$ (14)
While $M_{z}$ peaks at the intrinsic magnetic-dipole-like resonance of the inclusions, the pole of $\mu_{r}$ occurs when the bulk field in the denominator, $H_{\text{bulk}}=\langle H_{z}\rangle-M_{z}$, vanishes. The relative phase and dispersive behavior of $\langle H_{z}\rangle$ therefore shift the divergence of $\mu_{r}$ to longer wavelengths than the maximum of $M_{z}$.
1.3 3.2 Excitation of Transverse Magnetic (TM) Resonances
To excite the TM resonances of the doped BR cavity, it is illuminated by a normally incident plane wave with its electric field polarized along the $z$-axis. The transmittance and field profile of the TM1-like mode are shown in Fig. 9. For a period of $p=500~\mathrm{nm}$, Fig. 9a demonstrates that only the TM1-like mode exists, which is confined inside the NCs. In contrast, Fig. 9b, corresponding to $p=750~\mathrm{nm}$, reveals the emergence of an additional mode at a shorter wavelength, primarily confined in the regions between the NCs. When the period is further increased to $p=900~\mathrm{nm}$, both modes undergo a red shift, exhibit an enhanced $Q$-factor, and become spectrally closer to each other. At the same time, other modes also appear at shorter wavelengths, with their fields predominantly distributed in the Bragg mirror layers rather than within the cavity core.
Figure 9: Transmittance and the resonant-mode field distribution of a BR cavity with the same parameters as in Fig. 5, loaded with an array of cylindrical NPs of radius $R=150~\mathrm{nm}$, for periods of (a) $p=500~\mathrm{nm}$; (b) $p=750~\mathrm{nm}$; and (c) $p=900~\mathrm{nm}$. All structures are illuminated by a normally incident plane wave (along $y$-axis) with the electric field $E$ along $z$-axis (TM, $E_{z}$).
1.4 3.3 Tunability of the Hybrid Bragg-Mie Modes Beyond the Size Parameter
In addition to the conventional size parameters, $x_{i}=2\pi n_{i}r/\lambda$ and $x_{e}=2\pi n_{e}r/\lambda$, which depend on wavelength, particle size and the refractive indices of the particle and surrounding medium,13 the hybrid Bragg–Mie modes can be further tuned by external and structural parameters. Specifically, varying the angle of incidence and the bare-cavity cutoff wavelength introduces additional means for controlling the resonance wavelength of the Bragg-Mie modes. These parameters enable tunability of the hybrid resonances beyond what is achievable with the size parameter alone, offering a richer degree of freedom for engineering their spectral position and field confinement. For instance, this approach can enable straightforward fine tuning and alignment of the mode resonances with the magnetic dipole resonances of doped atoms or quantum dots, facilitating enhancement or selective excitation of their intrinsic magnetic dipole transitions.
Figure 10 illustrates this tunability for both the TE0-like (Fig. 10a) and TM1-like (Fig. 10b) modes by showing the transmittance of a doped BR cavity with the same parameters as in Fig. 5, for $p=500~\mathrm{nm}$ and different bare-cavity cutoff wavelengths of $\lambda_{c}=1450~\mathrm{nm}$, $\lambda_{c}=1500~\mathrm{nm}$, and $\lambda_{c}=1550~\mathrm{nm}$. In both cases, the resonance exhibits a blueshift for smaller cutoff wavelengths and a redshift for larger cutoff wavelengths.
Figure 10: Transmittance spectra of a doped BR cavity with $p=500~\mathrm{nm}$ and different cutoff wavelengths $\lambda_{c}=1450~\mathrm{nm}$ (red dashed), $\lambda_{c}=1500~\mathrm{nm}$ (blue solid), and $\lambda_{c}=1550~\mathrm{nm}$ (green dotted). (a) TE0-like mode, (b) TM1-like mode.
Figure 11 shows the TE-polarized transmittance of the structure in Fig. 10 for increasing AoI. The lower- and higher-energy modes shift and move closer together, illustrating an angle-tunable response; the lower-energy branch blue-shifts across the range, while the higher-energy branch shows a weaker, non-monotonic shift at larger angles.
Figure 11: TE-polarized transmittance of a structure with the same parameters of Fig. 10, and a cutoff wavelength of $1500~\mathrm{nm}$. Results are shown for incidence angles of $0^{\circ}$ (solid curve), $25^{\circ}$ (dashed curve), and $50^{\circ}$ (dash-dotted curve).
Figure 12 shows the TM-polarized transmittance of the same structure as in Fig. 11, but with $p=750~\mathrm{nm}$ and AoIs of $0^{\circ}$, $5^{\circ}$, and $15^{\circ}$. As seen in Fig. 12a, for nonzero AoIs a new resonance emerges with a markedly higher $Q$-factor, particularly at smaller angles (e.g., $Q\approx 1.5\times 10^{4}$ for AoI $5^{\circ}$, compared to $Q\approx 700$ for the TM1-like mode). This resonance is structurally similar to the TM1-like mode but exhibits a $90^{\circ}$ rotated field distribution. With increasing AoI, the new mode undergoes a red shift, while the TM1-like mode gradually vanishes. As illustrated in Fig. 12b for an AoI of $15^{\circ}$, additional resonances also appear at shorter wavelengths, with their fields predominantly distributed in the Bragg mirror layers rather than in the cavity core.
Figure 12: TM-polarized transmittance of a BR cavity with $N=10$ pairs of cladding layers on each side of a half-wave core, designed for a cutoff wavelength of $1500~\mathrm{nm}$, and loaded with an array of cylindrical NPs of radius $R=150~\mathrm{nm}$ and period $p=750~\mathrm{nm}$. (a) Results for incidence angles of $0^{\circ}$ (dashed curve) and $5^{\circ}$ (solid curve). (b) Results for an incidence angle of $15^{\circ}$.
The modes discussed above, including the TE- and TM-like resonances confined inside the NCs, the modes localized between the NCs, and the new high-$Q$ resonance emerging at nonzero AoI in Fig. 12a, can all be identified as high-FoM NZI modes of the doped BR cavity. This is confirmed by the dispersion curves shown in Fig. 13a in the wavelength range of 1300-1650 nm, together with their corresponding transmission peaks presented in Fig. 13b.
Figure 13: (a) Real and imaginary parts of the effective refractive index of TE and TM modes, together with their corresponding field distributions, in a BR cavity with $N=14$ pairs of cladding layers on each side of a half-wave core, loaded with an array of cylindrical NPs of radius $R=150~\mathrm{nm}$ and refractive index $n=3.5$, with a period of $p=750~\mathrm{nm}$. (b) Transmission spectrum of the same cavity for TE (solid curve) and TM (dashed/dotted curves) modes.
3.4 Magnetic Purcell Factor Enhancement in NC-Doped ENZ Cavity
To quantify the spontaneous-emission enhancement in the NC-doped ENZ Bragg cavity, we evaluate the PF and compare it with that of a single NC or an array of NCs embedded in an infinite homogeneous medium. The formulation follows the Green-tensor approach previously introduced for two-dimensional (2D) photonic systems 37, extended here to account for both electric and magnetic dipole emitters by means of the corresponding electric and magnetic Green functions.
For an electric dipole $\mathbf{p}=p_{0}\,\mathbf{u}_{p}$, with orientation $\mathbf{u}_{p}$ at position $\mathbf{r}_{0}$, the normalized decay rate is given by the LDOS expression 38
$\frac{\Gamma(\omega)}{\Gamma_{0}(\omega)}=\frac{6\pi}{k}\,\Im\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{p}\right\},\qquad k=\omega/c.$ (15)
Using the relation between the Green tensor and the local electric field (Appendix C), one obtains the relation 2
$\frac{\Gamma(\omega)}{\Gamma_{0}(\omega)}=-\,\frac{6\pi c}{\mu_{0}\,\omega^{3}\,|p_{0}|}\,\Re\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{E}(\mathbf{r}_{0},\omega)\right\}.$} (16)
A parallel derivation applies to a magnetic dipole $\mathbf{m}=m_{0}\,\mathbf{u}_{m}$, with orientation $\mathbf{u}_{m}$ at $\mathbf{r}_{0}$. Starting from the magnetic LDOS and invoking the relation between $\mathbf{H}$ and the magnetic Green tensor (Appendix C), we obtain
[…]
For our 2D cavity, the PF is evaluated by exciting the structure with either an electric or a virtual magnetic line current placed at the center of an NC in the core, oriented along $z$. The PF is then obtained from the above expressions by computing the real part of the corresponding $z$-component of the field at the source position and normalizing to its free-space value.
Figure 14: (a) Real and imaginary parts of the effective refractive index of TE and TM modes, together with their corresponding field distributions, in a BR cavity with $N=14$ pairs of cladding layers on each side of a half-wave core, loaded with an array of cylindrical NCs of radius $R=150~\mathrm{nm}$ and refractive index $n=3.5$, with a period of $p=750~\mathrm{nm}$. (b) Transmission spectrum of the same cavity for TE (solid curve) and TM (dashed curve) modes.
Figure 14a compares the calculated PF spectra at the center of a single NC in free space (dashed line), an array of 25 NCs in free space, and the same NC array embedded in the Bragg cavity (solid line). The embedded array exhibits a nearly 35-fold Purcell enhancement relative to the isolated NC or free-space array. The corresponding electric and magnetic field distributions at resonance are shown in Figs. 14b and 14c. In these simulations, the Bragg cavity consists of $N=14$ cladding-layer pairs on each side of a half-wave core, loaded with an array of 25 NCs of radius $R=150~\mathrm{nm}$ and refractive index $n=3.5$, with period $p=750~\mathrm{nm}$. The cavity width is $W=18.75~\mu\mathrm{m}$.
The magnetic hot-spot formation and Purcell enhancement effects discussed so far for arrays of particles are not limited to circular-cylinder geometries. As illustrated in Fig. 15, replacing the circular cross-section NCs of Fig. 14 with square cross-section NCs yields similar magnetic hot spots and even stronger PF values, exceeding $10^{3}$. The side length of each square NC is $300~\mathrm{nm}$, while all other structural parameters are kept identical to those in Fig. 14.
Figure 15: Magnetic Purcell factor spectrum for an array of square cross-section NCs with side length $300~\mathrm{nm}$ embedded in the Bragg cavity, with all other structural parameters identical to Fig. 14. A pronounced resonance is observed near $1505~\mathrm{nm}$, where the PF exceeds $10^{3}$. The insets show the corresponding electric ($|E|$) and magnetic ($|H|$) field distributions at resonance, highlighting strong magnetic hot-spot confinement inside the NCs.
The Purcell enhancement effect can also be observed in higher-order NZI modes of the cavity, which arise from reflections at the cavity sidewalls due to imperfect impedance matching with the surrounding medium, thereby forming standing-wave patterns inside the cavity. Interestingly, by tuning the cavity width, the PF at the center of cavity, associated with these higher-order modes can be optimized. For example, as shown in Fig. 16a, for an array of 25 NCs ($R=150~\mathrm{nm}$ and $p=500~\mathrm{nm}$) embedded in a Bragg cavity with $N=14$ and a width of $W=12.9~\mu\mathrm{m}$, the PF reaches values as high as 4600 for a higher-order NZI TE0-like mode, with the corresponding magnetic field profile shown in Fig. 16c.
Figure 16: Purcell enhancement in higher-order NZI modes of a Bragg cavity with $N=14$ cladding-layer pairs and width $W=12.9~\mu\mathrm{m}$, loaded with an array of 25 NCs of radius $R=150~\mathrm{nm}$ and period $p=500~\mathrm{nm}$. (a) Magnetic Purcell factor spectrum showing multiple resonances, with a maximum PF $\sim 4600$ at $\lambda=1356.4~\mathrm{nm}$. (b) Magnetic field distribution $\Re(H_{z})$ at $\lambda=1372~\mathrm{nm}$ corresponding to a lowest-order cavity mode. (c) Magnetic field distribution $\Re(H_{z})$ at $\lambda=1356.4~\mathrm{nm}$, corresponding to a highest-order NZI TE0-like mode responsible for the giant PF enhancement.
The electric Purcell enhancement in NC-doped Bragg cavities can be investigated for different orders of NZI TM-polarized modes by exciting the structure with electric line currents placed at the electric hot spots of the modes. It should be noted that the use of electric or virtual magnetic line currents is purely a computational tool to excite the corresponding TM or TE modes; in practice, the TE- and TM-polarized modes can be excited by plane-wave illumination with the appropriate polarization. Figure 17 presents the corresponding calculations for the same cavity configuration as in Fig. 16, but with a reduced width of $W=12.5~\mu\mathrm{m}$ and with electric rather than magnetic line-current excitation. The resulting electric PF spectrum, shown in Fig. 17a, exhibits a series of sharp resonances corresponding to different orders of NZI TM-like modes. The field distributions for the lowest-order mode at $\lambda=1399.8~\mathrm{nm}$ are shown in Figs. 17b and 17c, illustrating the magnitudes of the electric field ($|E|$) and magnetic field ($|H|$), respectively. Both profiles demonstrate strong field localization within the NCs, with the inset enlargements highlighting the distinct electric hot spots characteristic of the resonant TM mode.
Figure 17: Electric Purcell enhancement in a Bragg cavity with $N=14$ cladding-layer pairs and width $W=12.5~\mu\mathrm{m}$, loaded with an array of 25 NCs of radius $R=150~\mathrm{nm}$ and period $p=500~\mathrm{nm}$. (a) Electric Purcell factor spectrum (b) Electric field distribution $|E|$ and (c) magnetic field distribution $|H|$ for the lowest-order TM-like mode at $\lambda=1399.8~\mathrm{nm}$. Insets highlight the strong field localization within the NCs and the distinct electric hot spots responsible for the observed enhancement.
4. Concluding Remarks
We have demonstrated that photonic doping of ENZ Bragg cavities with arrays of dielectric NPs provides a versatile platform for engineering high-$Q$ NZI modes with strong electric and magnetic responses. By combining the dispersive properties of near-cutoff Bragg cavities with the resonant character of embedded NPs, hybrid Bragg–Mie modes emerge that can be tuned not only through the conventional size parameter, but also via structural and external degrees of freedom such as cavity width, cutoff wavelength, and angle of incidence. These modes support both electric and magnetic hot spots with unprecedented confinement, leading to PF enhancements of more than two orders of magnitude beyond those of isolated NPs. Importantly, the approach is not limited to circular geometries, as demonstrated for square cross-section NPs, and extends naturally to higher-order cavity modes supported by sidewall reflections. The resulting isolated, spectrally pure resonances with dipolar or multipolar character realize optical-frequency analogues of PMC and EMNZ behavior, enabled by the emergence of magnetic and electric hot spots. Our findings establish NC-doped ENZ Bragg cavities as a powerful paradigm for achieving ultra-narrow-bandwidth resonances with extreme field localization, offering new opportunities for low-threshold nonlinear optics, magnetic dipole spectroscopy, and quantum light–matter interactions.
Appendix A
We consider the Mie coefficient corresponding to TE modes of an infinite lossless cylinder:
$a_{n}=\frac{x_{i}J_{n}(x_{i})J_{n}^{\prime}(x_{e})-x_{e}J_{n}(x_{e})J_{n}^{\prime}(x_{i})}{x_{i}J_{n}(x_{i}){H_{n}^{(1)}}^{\prime}(x_{e})-x_{e}H_{n}^{(1)}(x_{e})J_{n}^{\prime}(x_{i})}$ (18)
The extinction coefficient is proportional to the real part of $a_{n}$, which can be written in terms of the real nominator (N) and complex denominator (D) of the coefficient as:
$\Re(a_{n})\approx\Re\left(\frac{1}{D}\right)\cdot N=\frac{\Re(D)\cdot N}{\Re(D)^{2}+\Im(D)^{2}}$ (19)
where $\Re(D)$ and $\Im(D)$ are the real and imaginary parts of $D$. We show that in the ENZ regime, the real part of the Mie coefficient $a_{n}$ (or $b_{n}$) behaves like a Lorentzian function normalized to unity at resonance:
$L(\omega)=\frac{\gamma^{2}}{\Delta\omega^{2}+\gamma^{2}}.$ (20)
In this representation, $\Im(D)$ vanishes at the resonance frequency and thus representing the role of the resonance detuning $\Delta\omega$, while the $\Re(D)$ approaches zero as the surrounding refractive index tends to zero, corresponding to the narrowing Lorentzian linewidth $\gamma$.
First, for $n=0$, using the small-argument approximations:
$H_{0}^{(1)}(x_{e})$ $\approx\left(1-\frac{x_{e}^{2}}{4}\right)+i\cdot\frac{2}{\pi}\left(\ln\left(\frac{x_{e}}{2}\right)+\gamma\right),$ (21)
$H_{0}^{(1)\prime}(x_{e})$ $\approx-\frac{x_{e}}{2}+i\cdot\left(-\frac{2}{\pi x_{e}}\right)$ (22)
the real and imaginary components of the denominator are obtained as:
$\Re(D)$ $=-\frac{x_{i}x_{e}}{2}J_{0}(x_{i})-x_{e}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right),$ (23)
$\Im(D)$ $=-\frac{2x_{i}J_{0}(x_{i})}{\pi x_{e}}-\frac{2x_{e}J_{0}^{\prime}(x_{i})}{\pi}\left(\ln\left(\frac{x_{e}}{2}\right)+\gamma\right).$ (24)
And using the small-argument approximations for the Bessel function:
$J_{0}(x_{e})$ $\approx 1-\frac{x_{e}^{2}}{4},$ (25)
$J_{0}^{\prime}(x_{e})$ $\approx-\frac{x_{e}}{2},$ (26)
the simplified expression for the numerator becomes:
$N\approx-\frac{1}{2}x_{i}x_{e}J_{0}(x_{i})-x_{e}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right).$ (27)
Therefore, when $x_{e}\to 0$, the resonance condition is obtained from Eq. (24) as $J_{0}(x_{i})=0$, and for the resonance width we have:
$\gamma_{0}\propto-x_{e}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right),$ (28)
which tends to zero as $x_{e}\to 0$.
For $n>0$, using the small-argument approximations:
$H_{n}^{(1)}(x_{e})\approx\frac{1}{n!}\left(\frac{x_{e}}{2}\right)^{n}-\frac{1}{(n+1)!}\left(\frac{x_{e}}{2}\right)^{n+2}+i\cdot\frac{(n-1)!}{\pi}\left(\frac{2}{x_{e}}\right)^{n}$ (29)
${H_{n}^{(1)}}^{\prime}(x_{e})\approx\frac{n}{2n!}\left(\frac{x_{e}}{2}\right)^{n-1}-\frac{(n+2)}{2(n+1)!}\left(\frac{x_{e}}{2}\right)^{n+1}-i\cdot\frac{n(n-1)!}{2\pi}\left(\frac{2}{x_{e}}\right)^{n+1}$ (30)
the real and imaginary parts of the denominator are given by:
$\Re(D)=$ $\frac{1}{2n!}\bigg[nx_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n-1}$ (31)
$\quad+\frac{1}{2(n+1)!}\bigg[-(n+2)x_{i}J_{n}(x_{i})+x_{o}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n+1}$
$\Im(D)=\frac{(n-1)!}{2\pi}\bigg[nx_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{2}{x_{e}}\right)^{n+1}$ (32)
Now we consider the numerator of the Mie coefficient $b_{n}$
$N(x_{e})=x_{i}J_{n}(x_{i})J_{n}^{\prime}(x_{e})-x_{e}J_{n}(x_{e})J_{n}^{\prime}(x_{i}),$
and expand the Bessel functions up to two terms:
$J_{n}(x_{e})\approx\frac{1}{n!}\left(\frac{x_{e}}{2}\right)^{n}-\frac{1}{(n+1)!}\left(\frac{x_{e}}{2}\right)^{n+2}$ (33)
$J_{n}^{\prime}(x_{e})\approx\frac{1}{2(n-1)!}\left(\frac{x_{e}}{2}\right)^{n-1}-\frac{1}{2(n+1)!}\left(\frac{x_{e}}{2}\right)^{n+1}$ (34)
to get:
$N$ $\approx\bigg[nx_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n-1}$ (35)
$-\frac{1}{2(n+1)!}\bigg[x_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n+1}.$
We see from Eq. 32 that the resonance condition is met by:
$nx_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})=0,$ (36)
which for $x_{e}\to 0$, reduces to:
$J_{n}(x_{i})=0.$ (37)
And for the resonance width we have:
$\gamma_{n}\propto\frac{-1}{2(n+1)!}\bigg[(n+2)x_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n+1},$ (38)
or:
$\gamma_{n}\propto\frac{1}{2(n+1)!}\bigg[x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n+1},$ (39)
which tends to zero as $x_{e}\to 0$.
In the case where the electric field is polarized parallel to the cylinder axis (TM polarization), the extinction efficiency is again proportional to the real part of the scattering coefficients $b_{n}$, expressed as:
$b_{n}=\frac{x_{e}J_{n}(x_{i})J_{n}^{\prime}(x_{e})-x_{i}J_{n}(x_{e})J_{n}^{\prime}(x_{i})}{x_{e}J_{n}(x_{i}){H_{n}^{(1)}}^{\prime}(x_{e})-x_{i}H_{n}^{(1)}(x_{e})J_{n}^{\prime}(x_{i})}.$ (40)
Using the small-argument expansions for $n=0$:
$H_{0}^{(1)}(x_{e})$ $\approx\left(1-\frac{x_{e}^{2}}{4}\right)+i\cdot\frac{2}{\pi}\left(\ln\left(\frac{x_{e}}{2}\right)+\gamma\right),$ (41)
$H_{0}^{(1)\prime}(x_{e})$ $\approx-\frac{x_{e}}{2}+i\cdot\left(-\frac{2}{\pi x_{e}}\right),$ (42)
$J_{0}(x_{e})$ $\approx 1-\frac{x_{e}^{2}}{4},$ (43)
$J_{0}^{\prime}(x_{e})$ $\approx-\frac{x_{e}}{2},$ (44)
the real and imaginary parts of the denominator become:
$\Re(D)$ $=-\frac{x_{e}^{2}}{2}J_{0}(x_{i})-x_{i}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right),$ (45)
$\Im(D)$ $=-\frac{2}{\pi}J_{0}(x_{i})-\frac{2x_{i}}{\pi}J_{0}^{\prime}(x_{i})\left(\ln\left(\frac{x_{e}}{2}\right)+\gamma\right)$ (46)
and the numerator becomes:
$N\approx-\frac{1}{2}x_{e}^{2}J_{0}(x_{i})-x_{i}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right)$ (47)
From Eq. (45) and (46) we see that to remove the divergence of the denominator, we should have:
$J_{0}^{\prime}(x_{i})=0,$ (48)
and consequently, $\Re(D)$ vanishes when $x_{e}\to 0$. However, $\Im(D)$ remains non-vanishing in the small-size-parameter limit, preventing the emergence of a resonance condition. So, as it can also be inferred from Fig. 2b, the TM0 mode of the infinite cylinder does not support a true resonance.
For $n>0$ we use the small-argument approximations:
$H_{n}^{(1)}(x_{e})\approx\frac{1}{n!}\left(\frac{x_{e}}{2}\right)^{n}+i\cdot\frac{(n-1)!}{\pi}\left(\frac{2}{x_{e}}\right)^{n}$ (49)
${H_{n}^{(1)}}^{\prime}(x_{e})\approx\frac{n}{2n!}\left(\frac{x_{e}}{2}\right)^{n-1}-i\cdot\frac{n(n-1)!}{2\pi}\left(\frac{2}{x_{e}}\right)^{n+1},$ (50)
to obtain the real and imaginary parts of the denominator as:
$\Re(D_{n})=\left(\frac{x_{e}}{2}\right)^{n}\cdot\frac{1}{n!}\left[\frac{n}{2}J_{n}(x_{i})-x_{i}J_{n}^{\prime}(x_{i})\right]+\left(\frac{x_{e}}{2}\right)^{n+2}\left[\frac{x_{i}J_{n}^{\prime}(x_{i})}{(n+1)!}-\frac{(n+2)}{2(n+1)!}J_{n}(x_{i})\right]$ (51)
$\Im(D_{n})=-\left(\frac{2}{x_{e}}\right)^{n}\cdot\frac{(n-1)!}{\pi}\left[\frac{n}{2}J_{n}(x_{i})+x_{i}J_{n}^{\prime}(x_{i})\right]$ (52)
Using only the leading-order terms of the Bessel functions:
$J_{n}(x_{e})\approx\frac{1}{n!}\left(\frac{x_{e}}{2}\right)^{n},\quad J_{n}^{\prime}(x_{e})\approx\frac{1}{2(n-1)!}\left(\frac{x_{e}}{2}\right)^{n-1}$ (53)
the numerator becomes:
$N\approx\left(\frac{x_{e}}{2}\right)^{n}\cdot\frac{1}{n!}\left[\frac{n}{2}J_{n}(x_{i})-x_{i}J_{n}^{\prime}(x_{i})\right]$ (54)
so, from Eqs. (32) and (33) we see that the resonance condition is satisfied when:
$\frac{n}{2}J_{n}(x_{i})+x_{i}J_{n}^{\prime}(x_{i})=0$ (55)
and for the resonance width we have:
$\gamma_{n}\propto\Re(D),$ (56)
indicating that the resonance width tends to zero as $x_{e}\to 0$.
Appendix B
Mie resonance of spherical NPs in ENZ media
To analyze the ENZ limit $x_{e}\to 0$, we use the small-argument expansions of the functions:
$\psi_{n}(x_{e})=\frac{x_{e}^{n+1}}{(2n+1)!!}+\mathcal{O}(x_{e}^{n+3}),\quad\psi_{n}^{\prime}(x_{e})=\frac{(n+1)x_{e}^{n}}{(2n+1)!!}+\mathcal{O}(x_{e}^{n+2})$ (57)
$\xi_{n}(x_{e})=\frac{x_{e}^{n+1}}{(2n+1)!!}+i\left(-\frac{(2n-1)!!}{x_{e}^{n}}\right)+\mathcal{O}(x_{e}^{n+3})+i\mathcal{O}(x_{e}^{2-n})$ (58)
$\xi_{n}^{\prime}(x_{e})=\frac{(n+1)x_{e}^{n}}{(2n+1)!!}+i\left(n\frac{(2n-1)!!}{x_{e}^{n+1}}\right)+\mathcal{O}(x_{e}^{n+2})+i\mathcal{O}(x_{e}^{1-n})$ (59)
and write the real part of $b_{n}$ in terms of its nominator ($N$) and denominator ($D$) as:
$\Re(a_{n})\approx\Re\left(\frac{1}{D}\right)\cdot N=\frac{\Re(D)\cdot N}{\Re(D)^{2}+\Im(D)^{2}}$ (60)
Then the numerator $N$, $\Re(D)$ and $\Im(D)$ are calculated as:
$N=\frac{x_{e}^{n+1}}{(2n+1)!!}\left((n+1)\psi_{n}(x_{i})-x_{i}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{n+3}),$ (61)
$\Re(D)=\frac{x_{e}^{n+1}}{(2n+1)!!}\left((n+1)\psi_{n}(x_{i})-x_{i}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{n+3}),$ (62)
$\Im(D)=\frac{(2n-1)!!}{x_{e}^{n}}\left(n\psi_{n}(x_{i})+x_{i}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{2-n}).$ (63)
From these results we can see that the dipolar or multipolar magnetic resonances occur when $\Im(D)=0$ or:
$n\psi_{n}(x_{i})+x_{i}\psi_{n}^{\prime}(x_{i})=0,$ (64)
which is consistent with the blue curves in Fig. 3b. For the resonance widths we get:
$\gamma_{n}\propto\frac{x_{e}^{n+1}}{(2n+1)!!}\left((n+1)\psi_{n}(x_{i})-x_{i}\psi_{n}^{\prime}(x_{i})\right)$ (65)
tending to zero as $x_{e}\to 0$.
Similarly, for the $a_{n}$ coefficients we obtain:
$N=\frac{x_{e}^{n}}{(2n+1)!!}\left((n+1)x_{i}\psi_{n}(x_{i})-x_{e}^{2}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{n+2}),$ (66)
$\Re(D)=\frac{x_{e}^{n}}{(2n+1)!!}\left((n+1)x_{i}\psi_{n}(x_{i})-x_{e}^{2}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{n+2}),$ (67)
$\Im(D)=(2n-1)!!\left(\psi_{n}(x_{i})n\frac{x_{i}}{x_{e}^{n+1}}+\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{1-n}),$ (68)
indicating that in the ENZ regime ($x_{e}\to 0$), the resonance width, proportional to $\Re(D)$, approaches zero. Although the divergence of the first term in $\Im(D)$ suggests that resonances are suppressed in this regime, a more careful analysis which follows, reveals that for any arbitrarily small $x_{e}$, there always exist roots of $\Im(D)$ given by
$x_{i}=\frac{\alpha}{2}+\frac{1}{2n}\sqrt{n^{2}\alpha^{2}-4nx_{e}^{n+1}},$ (69)
located near the zeros $\alpha$ of the function $\psi_{n}(x)$, which is consistent with the red curves in Fig. 3b.
The roots of $\Im(D)$
Let $\alpha$ denote a simple zero of $\psi_{n}(x_{i})$, that is,
$\psi_{n}(\alpha)=0,\quad\text{and}\quad\psi_{n}^{\prime}(\alpha)\neq 0.$ (70)
Expanding $\psi_{n}(x_{i})$ near $\alpha$ gives
$\psi_{n}(x_{i})\approx\psi_{n}^{\prime}(\alpha)(x_{i}-\alpha),$ (71)
and also,
$\psi_{n}^{\prime}(x_{i})\approx\psi_{n}^{\prime}(\alpha).$ (72)
Substituting into the function
$f(x_{i})=\psi_{n}(x_{i})n\frac{x_{i}}{x_{e}^{n+1}}+\psi_{n}^{\prime}(x_{i}),$ (73)
we find
$f(x_{i})\approx\psi_{n}^{\prime}(\alpha)\left(n\frac{x_{i}(x_{i}-\alpha)}{x_{e}^{n+1}}+1\right).$ (74)
Setting $f(x_{i})=0$, we obtain
$nx_{i}(x_{i}-\alpha)+x_{e}^{n+1}=0,$ (75)
or equivalently,
$nx_{i}^{2}-n\alpha x_{i}+x_{e}^{n+1}=0.$ (76)
which has the solutions:
$x_{i}=\frac{\alpha}{2}\pm\frac{1}{2n}\sqrt{n^{2}\alpha^{2}-4nx_{e}^{n+1}}.$ (77)
As $x_{e}\to 0$, the square root behaves like
$\sqrt{n^{2}\alpha^{2}-4nx_{e}^{n+1}}\approx n\alpha,$ (78)
thus the ++ root approaches $x_{i}=\alpha$ while the - root approaches $x_{i}=0$. Since we seek a root near $\alpha$, we select the positive sign root:
$x_{i}=\frac{\alpha}{2}+\frac{1}{2n}\sqrt{n^{2}\alpha^{2}-4nx_{e}^{n+1}}.$} (79)
Appendix C:
.5 Electric and Magnetic Dipoles, Green Tensor, LDOS, and Purcell Factor
C.1. Conventions
We use time-harmonic fields with the phasor convention $e^{-i\omega t}$. Free-space constants are $\varepsilon_{0}$ and $\mu_{0}$, with $c=1/\sqrt{\varepsilon_{0}\mu_{0}}$ and $k_{0}=\omega/c$. Throughout, material permeabilities are taken as $\mu(\mathbf{r},\omega)=\mu_{0}$.
C.2. Electric dipole: field and Purcell factor
Applying the inhomogeneous vector Helmholtz equation
$\nabla\times\nabla\times\mathbf{E}(\mathbf{r})-k_{0}^{2}\,\varepsilon(\mathbf{r},\omega)\,\mathbf{E}(\mathbf{r})=i\omega\mu_{0}\,\mathbf{J}(\mathbf{r}),$ (80)
the electric dyadic Green tensor $\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)$ is defined as the solution of
$\nabla\times\nabla\times\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)-k_{0}^{2}\,\varepsilon(\mathbf{r},\omega)\,\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)=\mathbf{I}\,\delta(\mathbf{r}-\mathbf{r}^{\prime}),$ (81)
satisfying the outgoing-wave radiation condition. The total field produced by an arbitrary current distribution $\mathbf{J}(\mathbf{r}^{\prime})$ is obtained as
$\mathbf{E}(\mathbf{r})=i\omega\mu_{0}\int\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)\,\mathbf{J}(\mathbf{r}^{\prime})\,\mathrm{d}^{3}r^{\prime}.$ (82)
A point electric dipole $\mathbf{p}$ at $\mathbf{r}_{0}$ is represented by a polarization $\mathbf{P}(\mathbf{r})=\mathbf{p}\,\delta(\mathbf{r}-\mathbf{r}_{0})$, with bound current
$\mathbf{J}(\mathbf{r})=-\,i\omega\,\mathbf{P}(\mathbf{r})=-\,i\omega\,\mathbf{p}\,\delta(\mathbf{r}-\mathbf{r}_{0}).$ (83)
Substituting (83) into (82) gives
$\mathbf{E}(\mathbf{r})=\mu_{0}\omega^{2}\,\mathbf{G}(\mathbf{r},\mathbf{r}_{0};\omega)\,\mathbf{p},\qquad\mathbf{E}(\mathbf{r}_{0})=\mu_{0}\omega^{2}\,\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\,\mathbf{p}.$ (84)
We also write $\mathbf{p}=p_{0}\,\mathbf{u}_{p}$ with $|\mathbf{u}_{p}|=1$.
The normalized electric decay rate (Purcell factor) for a dipole oriented along $\mathbf{u}_{p}$ at $\mathbf{r}_{0}$ is
$\frac{\Gamma(\omega)}{\Gamma_{0}(\omega)}=\frac{6\pi}{k}\,\Im\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{p}\right\},\qquad k=\frac{\omega}{c}.$ (85)
From (84),
$\mathbf{u}_{p}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{p}=\frac{\mathbf{u}_{p}\!\cdot\!\mathbf{E}(\mathbf{r}_{0})}{\mu_{0}\omega^{2}\,p_{0}}.$ (86)
Choosing the (arbitrary) source phase $p_{0}=i|p_{0}|$ and using $\Im\{X/i\}=-\Re\{X\}$ for any complex scalar $X$, we obtain
$\Im\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{G}\!\cdot\!\mathbf{u}_{p}\right\}=-\,\frac{\Re\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{E}(\mathbf{r}_{0})\right\}}{\mu_{0}\omega^{2}|p_{0}|}.$ (87)
Substituting (87) into (85) yields
$\frac{\Gamma(\omega)}{\Gamma_{0}(\omega)}=-\,\frac{6\pi c}{\mu_{0}\,\omega^{3}\,|p_{0}|}\,\Re\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{E}(\mathbf{r}_{0},\omega)\right\}.$} (88)
C.3. Magnetic dipole: field and Purcell factor
We include a fictitious magnetic current density $\mathbf{M}$ to treat magnetic dipoles on the same footing as electric ones (phasor convention $e^{-i\omega t}$):
$\nabla\times\mathbf{E}$ $=i\omega\mu_{0}\,\mathbf{H}\;-\;\mathbf{M},$ (89)
$\nabla\times\mathbf{H}$ $=\mathbf{J}\;-\;i\omega\varepsilon_{0}\varepsilon(\mathbf{r},\omega)\,\mathbf{E}.$ (90)
Let $\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)$ be the electric dyadic Green tensor defined in (81). One may verify by direct substitution in (89)–(90) that the magnetic field generated by an arbitrary $\mathbf{M}$ is
$\mathbf{H}(\mathbf{r})=i\omega\,\varepsilon_{0}\varepsilon(\mathbf{r},\omega)\int\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)\,\mathbf{M}(\mathbf{r}^{\prime})\,\mathrm{d}^{3}r^{\prime}.$ (91)
A point magnetic dipole $\mathbf{m}=m_{0}\,\mathbf{u}_{m}$ with $|\mathbf{u}_{m}|=1$, located at $\mathbf{r}_{0}$ is represented by
$\mathbf{M}(\mathbf{r})=-\,i\omega\,\mathbf{m}\,\delta(\mathbf{r}-\mathbf{r}_{0}),$ (92)
so that, evaluating (91) at the source point and denoting the local (host) permittivity by $\varepsilon_{b}(\omega)=\varepsilon(\mathbf{r}_{0},\omega)$,
$\mathbf{H}(\mathbf{r})=\varepsilon_{0}\varepsilon_{b}\,\omega^{2}\,\mathbf{G}(\mathbf{r},\mathbf{r}_{0};\omega)\,\mathbf{m},\qquad\mathbf{H}(\mathbf{r}_{0})=\varepsilon_{0}\varepsilon_{b}\,\omega^{2}\,\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\,\mathbf{m}.$ (93)
Let $c_{b}=1/\sqrt{\varepsilon_{0}\varepsilon_{b}\mu_{0}}$ be the wave speed and $k_{b}=\omega/c_{b}$ the wavenumber in the host at $\mathbf{r}_{0}$. The magnetic Purcell factor (normalized decay rate) is
$\frac{\Gamma_{m}(\omega)}{\Gamma_{m,0}(\omega)}=\frac{6\pi}{k_{b}}\,\Im\!\left\{\mathbf{u}_{m}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{m}\right\}.$ (94)
Projecting (93) along $\mathbf{u}_{m}$ gives
$\mathbf{u}_{m}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{m}=\frac{\mathbf{u}_{m}\!\cdot\!\mathbf{H}(\mathbf{r}_{0})}{\varepsilon_{0}\varepsilon_{b}\,\omega^{2}\,m_{0}}.$ (95)
Choosing the dipole phase $m_{0}=i|m_{0}|$ (so that $\Im\{X/i\}=-\Re\{X\}$) yields
[…]